40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 68 of 348  Not logged in ELOG logo
ID Date Author Type Category Subject
  14120   Tue Jul 31 22:50:18 2018 aaronUpdateOMCOMC Expected Refl Signal

I learned a lot about lasers this week from Siegman. Here are some plots that show the expected reflectivity off of the OMC for various mode matching cases.

The main equation to know is 11.29 in Siegman, the total reflection coefficient going into the cavity:

R=r-\frac{t^2}{r}\frac{g(\omega)}{1-g(\omega)}

Where r is the mirror reflectivity (assumed all mirrors have the same reflectivity), t is the transmissivity, and g is the complex round-trip gain, eq 11.18

g(\omega)=r_1r_2(r_3...)e^{-i\phi}e^{-\alpha_0p}

The second exponential is the loss; in Siegman the \alpha_0 is some absorption coecfficient and p is the total round trip length, so the product is just the total loss in a round trip, which I take to be 4*the loss on a single optic (50ppm each). \phi is the total round trip phase accumulation, which is 2\pi*detuning(Hz)/FSR. The parameters for the cavity can be found on the wiki.

I've added the ipynb to my personal git, but I can put it elsewhere if there is somewhere more appropriate. I think this is all OK, but let me know if something is not quite right.

Attachment 1: omcRefl.pdf
omcRefl.pdf
  14119   Tue Jul 31 08:17:55 2018 SteveUpdateSUSTrillium interface box was fixed,reinstalled & working

 

 

Attachment 1: all_OK.png
all_OK.png
  14118   Mon Jul 30 18:19:03 2018 KojiUpdateSUSTrillium interface box was fixed and reinstalled

The trillium interface box was removed from the rack.

The problem was the incorrect use of an under-spec TVS (Transient Voltage Suppression) diodes (~ semiconductor fuse) for the protection circuit.
The TVS diodes we had had the breakdown voltages lower than the supplied voltages of +/-20V. This over-voltage eventually caused the catastrophic breakdown of one of the diodes.

I don't find any particular reason to have these diodes during the laboratory use of the interface. Therefore, I've removed the TVS diodes and left them unreplaced. The circuit was tested on the bench and returned to the rack. All the cables are hooked up, and now the BRLMs look as usual.


Details

- The board version was found to be D1000749-v2

- There was an obvious sign of burning or thermal history around the components D17 and D14. The solder of the D17 was so brittle that just a finger touch was enough to remove the component.

- These D components are TVS diodes (Transient Voltage Suppression Diodes) manufactured by Littelfuse Inc. It is sort of a surge/overvoltage protector to protect rest of the circuit to be exposed to excess voltage. The specified component for D17/D14 was 5.0SMMDJ20A with reverse standoff voltage (~operating voltage) of 20V and the breakdown voltage of 22.20V(min)~24.50V(max). However, the spec sheet told that the marking of the proper component must be "5BEW" rather than "DEM," which is visible on the component. Some search revealed that the used component was SMDJ15A, which has the breakdown voltage of 16.70V~18.50V. This spec is way too low compared to the supplied voltage of +/-20V.

Attachment 1: P_20180730_173134.jpg
P_20180730_173134.jpg
Attachment 2: P_20180730_180151.jpg
P_20180730_180151.jpg
  14117   Mon Jul 30 16:11:54 2018 gautamUpdateSUSTrillium interface box is broken

[koji, steve, gautam]

We debugged this in the following way:

  1. Disconnect all fuses in the terminal blocks coming from the +/- 20 VDC Sorensens.
  2. Check that they are indeed isolated using DMM.
  3. Test blocks of fuses in order to identify where the problem is happening (i.e. plug fuses in, turn up Sorensen voltage knobs, look for current overload). We did things in the following order:
    • MC suspensions
    • BS, PRM and SRM
    • ITMY
    • ITMX
    • Trillium interface box.
  4. Turns out that the Trillium box is the culprit.
  5. Confirmed that the problem is in the trillium interface box and not in the seismometer itself by unplugging all cables leading out of the interface box, and checking that the problem persists when the box is powered on.

So for now, the power cable to the box is disconnected on the back end. We have to pull it out and debug it at some point.

Apart from this, megatron was un-sshable so I had to hard reboot it, and restart the MCautolocker, FSSslowPy and nds2 processes on it. I also restarted the modbusIOC processes for the PSL channels on c1auxex (for which the physical Acromag units sit in 1X5 and hence were affected by our work), mainly so that the FSS_RMTEMP channel worked again. Now, IMC autolocker is working fine, arms are locked (we can recover TRX and TRY~1.0), and everything seems to be back to a nominal state. Phew.

  14115   Mon Jul 30 11:05:44 2018 gautamUpdateSUSIFO SUS wonky

When I came in this morning:

  • PMC was unlocked.
  • Seis BLRMS were off scale.
  • ITMX OSEM LEDs were dark on the CRT monitor even though Sat Box was plugged in.

Checking status of slow machines, it looked like c1sus, c1aux, and c1iscaux needed reboots, which I did. Still PMC would not lock. So I did a burtrestore, and then PMC was locked. But there seemed to be waaaaay to much motion of MCREFL, so I checked the suspension. The shadow sensor EPICS channels are reporting ~10,000 cts, while they used to be ~1000cts. No unusual red flags on the CDS side. Everything looked nominal when I briefly came in at 6:30pm PT yesterday, not sure if anything was done with the IFO last night.

Pending further investigation, I'm leaving all watchdogs shutdown and the PSL shutter closed.

A quick look at the Sorensens in 1X6 revealed that the +/- 20V DC power supplies were current overloaded (see Attachment #1). So I set those two units to zero until we figure out what's going on. Possibly something is shorted inside the ITMX satellite box and a fuse is blown somewhere. I'll look into it more once Steve is back.

Attachment 1: IMG_7102.JPG
IMG_7102.JPG
  14114   Sun Jul 29 23:15:34 2018 poojaUpdateCamerasDeveloping CNN

Aim: To develop a convolutional neural network that resolves mirror motion from video.

Input : Previous simulated video of beam spot motion in pitch by applying 4 sine  waves of frquencies 0.2, 0.4, 0.1, 0.3 Hz  and amplitude ratios to frame size to be 0.1, 0.04, 0.05, 0.08 where random uniform noise ranging 0.05 has been added to amplitudes and frequencies. This is divided into train (0.4), validation (0.1) and test (0.5).

Model topology:

  • Number of filters = 2
  • Kernel size = 2
  • Size of pooling windows = 2
  •                                        ----->         Dense layer of 4 nodes  ---->    Output layer of 1 node 

         Activation:                      selu                                                  linear

Batch size = 32, Number of epochs = 128, loss function = mean squared error

Optimizer: Nadam ( learning rate = 0.00001, beta_1 = 0.8, beta_2 = 0.85)

Plots of CNN output & applied signal given in Attachment 1. The variation in loss value with epochs given in Attachment 2.

This needs to be further analysed with increasing random uniform noise over the pixels and by training CNN on simulated data of varying ampltides and frequencies for sine waves.

Attachment 1: conv_nn_varying_freq_amp_1.pdf
conv_nn_varying_freq_amp_1.pdf
Attachment 2: conv_nn_varying_freq_amp_2.pdf
conv_nn_varying_freq_amp_2.pdf
  14113   Sun Jul 29 20:03:02 2018 ranaUpdatePEMSeismometer temp control

While Shruti is re-building Kira's heater circuit, I looked up how to do one of these (i.e. what does a real EE say about how to build a current source?):

It turns out that there is an Analog Devices application note (AN-968) about this (as there usually is once we get tired of playing around and try to look up the right answer).

I've linked to the note and attached the recommended schematic for high current applications. We'll go ahead as is, but we'll make a PCB according to this App Note for the v3 circuit.

 

Attachment 1: Screen_Shot_2018-07-29_at_8.00.27_PM.png
Screen_Shot_2018-07-29_at_8.00.27_PM.png
  14112   Sun Jul 29 00:59:54 2018 KojiUpdateElectronicsCharacterization of Transimpedance Amplifier

You have this measurement problem when the IF bandwidth is larger than the measurement frequency. I suspect the IF bandwidth is 30kHz.

  14111   Sat Jul 28 22:16:49 2018 John MartynUpdate Characterization of Transimpedance Amplifier

Kevin and I meaured the transfer function of the photodiode circuit using the Jenne laser and agilent in the 40m lab. The attached figures depict our measured transfer function over the modulation frequency ranges of 30kHz-30MHz and 1kHz-30MHz when the power of the laser was set to 69 and 95 μW. These plots indicate a clear roll off frequency around 300 kHz. In addition, the plots beginning at 1kHz display unstable behavior at frequencies below 30kHz. I am not sure why there is such a sharp change in the transfer function around 30kHz, but I suspect this to be due to an issue with the agilent or photodiode. 

Attachment 1: PD_TF1.pdf
PD_TF1.pdf
Attachment 2: PD_TF2.pdf
PD_TF2.pdf
Attachment 3: PD_and_TIA_Transfer_Function_Measurements.zip
  14110   Sat Jul 28 00:45:11 2018 terra, sandrineSummaryThermal CompensationHeater measurements overview

[Sandrine, Koji, Terra]

Summary: We completed multiple scans at different heating powers for the reflector set up, observing unique HOM peak shifts of tens of kHz. We also observed HOM5 shifts with the cylinder set up. Initial Lorentzian fittings of the magnitude give tens of Hz resolution. I summarize the main week's work below. 

Set-up

Heater set-up is described in several previous elogs, but attachments #1 and #2 show the full heater set-up and wiring/pinouts in and out of vacuum, since we're all intimately aware of how confusing in-vacuum pinouts can be. We are not using the Sorenson power supply (as described in 14071); we just have the BKPrecision power supply 1735 sitting next to the ETMY rack and are manually going out to turn on/off. 

We've continued to use the scan setup described in elog 14086, which is run using /users/annalisa/postVent/AGfast.py. Step by step notes for setting up the scan, running the scans, and processing the scans are attached in notes.txt.

Inducing/witnessing HOMs

The aux input beam was already clipped and on wednesday (after Trans was centered, 14093) we also clipped the output aux beam with razor blade (angled vertically and horizontally, elog 14103) before PDA255; we clipped ~1/3 of the output beam. Attachment #3 shows before and after clipping output, where orange 'cold' == unclipped, black 'mean' == clipped (all in cold state). Up to HOM5 is visible. 

Measurements

Below is a summary of the available scan data. We also have cold (0A) scans CAR-HOM5 and full FSR scans for most configurations. 

Elliptic Reflector
current[A] voltage[V] power[W] scans
0.4 2 0.8 CAR-HOM3(x1)
0.5 3.4 1.7 CAR-HOM3(x1)
0.6 5 3.0 CAR-HOM3(x1)
0.8 9.4(9.7) 7.5(7.8) CAR-HOM5(>x5)
0.9 12 10.8 CAR-HOM5(x4)
1.09 17 18.5 CAR-HOM3

 

 

 

 

 

 

 

Cylinder + Lenses
current[A] voltage[V] power[W] scans
0.9 15 13.5 CAR-HOM5(odds x4)

We tried the cylinder set-up again tonight for the first time since inital try and can see shifts of HOM5 - see attachment #5; we haven't looked in detail yet, but it looks like odd modes are more effected, suggesting the ring heat pattern is off centered from the beam axis. 

Scan data is saved in the following format: users/annalisa/postVent/scandata/{reflector,cylinder}/{parsed,unparsed}/{CAR,HOM1,HOM2,HOM3,HOM4,HOM5}{_datetime}{_parsed,_unparsed}.{txt,pdf}

Minimum heating

On 7/26 we increased the power to the elliptical reflector heater in steps to find the minimum heater power required to see frequency shifts with our measurement setup. Lowest we can resolve is a shift in HOM3 with 1.7W (0.5A/3.4V). According to Annalisa's measurements in elog 14050, this would be something like 30-60 mW radiated power hitting the test mass. We only looked at CAR - HOM3 for this investigation; data for scans at 0.4A, 0.5A, 0.6A is available as indicated above.

Lorentizian Fitting

The Lorentzian fitting was done using the equation a + b / sqrt(1+((x-c)/d*2), where a = constant background, b = peak height above background, c = peak frequency, d = full width at half max. 

The fitting is still being edited and optimized. We will crop the data to zoom in around the peak more.

The Lorentzian fit of the magnitude shows ~10Hz of resolution. (See attachment 6 for the carrier at 8A and attachment 7 for HOM 1 at 9A)

We're working on fitting the full complex data.

 

 

Attachment 1: heater_setup.jpg
heater_setup.jpg
Attachment 2: heater_wiring.jpg
heater_wiring.jpg
Attachment 3: notes.txt
Notes for running scans:
1. when first turning on Agilent, set initial stuff
    > cd /users/annalisa/postVent/20180718
    > AGmeasure TFAG4395Atemplate.yml
2. tweak arm alignment and offset PLL
    > sitemap (then IFO --> ALIGN and also PSL --> AUX)
    > to increase 
3. make sure X-arm is misagligned (hit '! Misalign' button for ITMX, ETMX) 
3. run scan
    > python AGfast.py startfreq stopfreq points
... 36 more lines ...
Attachment 4: FSR_clipped.pdf
FSR_clipped.pdf
Attachment 5: cylinderHOM5.pdf
cylinderHOM5.pdf
Attachment 6: pt8A_CAR.pdf
pt8A_CAR.pdf
Attachment 7: pt9A_HOM1.pdf
pt9A_HOM1.pdf
  14109   Fri Jul 27 17:16:14 2018 SandrineUpdateThermal CompensationCopied working scripts for mode spectroscopy into new directory (modeSpec)

The scripts: AGfast.py, make HDF5.py, plotSpec_marconi.py, and SandrineFitv3.py were copied into the new directory modeSpec.

The path is: /opt/rtcds/caltech/c1/scripts/modeSpec

These scripts can still be found under Annalisa's directory under postVent.

  14108   Fri Jul 27 10:48:57 2018 SteveUpdateSUSBS oplev window

Yesterday I inspected this BS oplev viewport. The heavy connector tube was shorting to table so It was moved back towards the chamber. The connection is air tight with kapton tape temporarly.

 The beam paths are well centered. The viewport is dusty on the inside.

The motivation was to improve the oplev noise.

Attachment 1: BSOw_.jpg
BSOw_.jpg
Attachment 2: dustInsideBSO.jpg
dustInsideBSO.jpg
  14107   Fri Jul 27 02:30:51 2018 gautamUpdateGeneralGlitchy MC

Kevin and I saw some weird IMC / PEM BLRMS behaviour today - see Attached screenshot. Not sure what was happening with the IMC, but MCtrans was oscillating at ~3Hz for a good 20 minutes or so. I just killed the lock, and restarted MCautolocker on megatron. There was a strange feature in the 3-10Hz BLRMS around that time as well. All seems back to normal now...

Attachment 1: 38.png
38.png
  14106   Thu Jul 26 15:11:18 2018 SteveUpdateGeneral Viewports & coating of 2001

New optical quality BK-7 windows in 2001 [4 substrates ] AR coated R<0.75 % for 630-1064nm " Azury BLue" broadband : TRX, TRY, ITMY-Oplev &  ITMX-Oplev viewports.

The BS-Oplev and PRM-Oplev 10" CF with 5.38" diameter view was coated the same way. The window here is Corning 7056 Borosilicate

5 more BK-7 substrates were coated R <0.1% of 1064 nm "Golden Orange" Their location: IMC-IN, IFO-REF and OMC   The next vent we have to confirm optical quality window locations.

All other conflat flange viewports are 7056 Kovar sealed .

Technical notes of 2001 40m upgrade can be seen at LIGO-T010115- 00- R  ....page 14

Attachment 1: BK7window_Coatings.PDF
BK7window_Coatings.PDF BK7window_Coatings.PDF
  14105   Thu Jul 26 01:52:01 2018 terraUpdateThermal Compensationheater work update

Just a quick update: over the past few days we've taken (at least) 5 scans around each peak [carrier - HOM3] at 9.4V/0.8A, 4 scans around [carrier - HOM5] at 12V/0.9A hot state with the reflector setup. We also have (at least) 5 scans of carrier - HOM5 in cold state. I attach a rough overview of the peak magnitude shifts in the first attachment. Analysis ongoing. All data stored in annalisa/postVent/{date}

Initial shifts just based on rought peak placement in the meantime:

            [9.4V/0.8A]   [12V/0.9A]

HOM1    10 kHz         20 kHz

HOM2    18 kHz         28 kHz

HOM3     30 kHz        40 kHz

HOM4     N/A             26 kHz

HOM5     N/A             35 kHz

I also attach the heating thermal transient from today (12V/0.9A) as seen by the opLevs. We see a shorter time constant for pitch, longer for yaw, preceeded by a dip in yaw. Similar behavior yesterday for slightly less heating, though less pronounced pre-dip. The heater is offcentered on the optic horizontally; likely this is part of the induced yaw. The spikey stuff i removed is from people walking around inside during the transient.

I've left the heater and LSC off for the night. Heater off at 2:07 am local time.

Please don't touch the oplevs; we're taking a cool down measurement.

Attachment 1: OpLev_thermal_drift.pdf
OpLev_thermal_drift.pdf
Attachment 2: hotColdAll.pdf
hotColdAll.pdf
  14104   Wed Jul 25 22:46:15 2018 gautamConfigurationComputersNDS access from outside

After this work, I've been having some trouble getting data with Python NDS. Eventually, I figured out that the nds connection request has to be pointed at '131.215.115.200' (the address of the NAT router which faces the outside world), port 31200 (it used to work with 'nds40.ligo.caltech.edu' or '131.215.115.189'). So the following snippet in python allows a connection to be opened. Offline access of frame data via NDS2 now seems possible.

import nds2
conn = nds2.connection('131.215.115.200',31200)
Quote:
 

So far, ssh (22), web services (30889), and elog (8081, 8080) were tested. We also need to test megatron NDS port forwarding and rsync for nodus, too.Finally I turned off the firewall rules of shorewall on nodus as it is no longer necessary.

  14103   Wed Jul 25 14:45:59 2018 SandrineSummaryThermal CompensationETM Y Table AUX read out

Attached is a photo of the set up of the ETM Y table showing the AUX read out set up. 

Currently, the flip mount sends the AUX to the PDA255. Terra inserted a razor blade so the PDA255 will witness more HOMs. The laser is also sent to the regular PD and the CCD.

Attachment 1: EY_table_.JPG
EY_table_.JPG
  14101   Tue Jul 24 09:47:51 2018 gautamUpdateCamerasDeveloping neural networks on simulated video

I was thinking a little more about the way we are training the network for the current topology - because the network has no recurrent layers, I guess it has no memory of past samples, and so it doesn't have any sense of the temporal axis. In fact, Keras by default shuffles the training data you give it randomly so the time ordering is lost. So the training amounts to requiring the network to identify the center of the Gaussian beam and output that. So in the training dataset, all we need is good (spatial) coverage of the area in which the spot is most likely to move? Or is the idea to develop some tools to generate video with spot motion close to that on the ETM in lock, so that we can use it with a network topology that has memory? 

Quote:

This looks like good progress. Instead of fixed sines or random noise, you should generate now a time series for the motion which is random noise but with a power spectrum similar to what we see for the ETM pitch motion in lock. You can use inverse FFT to get the time series from the open loop OL spectra (being careful about edge effects)

  14100   Tue Jul 24 06:11:50 2018 ranaUpdateCamerasDeveloping neural networks on simulated video

This looks like good progress. Instead of fixed sines or random noise, you should generate now a time series for the motion which is random noise but with a power spectrum similar to what we see for the ETM pitch motion in lock. You can use inverse FFT to get the time series from the open loop OL spectra (being careful about edge effects).

Quote:

Aim: To develop a neural network that resolves mirror motion from video.

  14098   Mon Jul 23 09:58:52 2018 SteveSummaryVACRGA scan at day 6

 

 

Attachment 1: pd81-560Hz-d6.png
pd81-560Hz-d6.png
  14097   Sun Jul 22 14:01:07 2018 poojaUpdateCamerasDeveloping neural networks on simulated video

Aim: To develop a neural network that resolves mirror motion from video.

Since error was high for the same input as in my previous elog http://nodus.ligo.caltech.edu:8080/40m/14089

I modified the network topology by tuning the number of nodes, layers and learning rate so that the model fitted the sum of 4 sine waves efficiently, saved weights of the final epoch and then in a different program, loaded saved weights & tested on simulated video that's produced by moving beam spot from the centre of image by sum of 4 sine waves whose frequencies and amplitudes change with time.

Input : Simulated video of beam spot motion in pitch by applying 4 sine  waves of frquencies 0.2, 0.4, 0.1, 0.3 Hz  and amplitude ratios to frame size to be 0.1, 0.04, 0.05, 0.08. This is divided into train (0.4), validation (0.1) and test (0.5).

Model topology:

                                          Input               -->                  Hidden layer               -->                    Output layer                                  

                                                                                          8 nodes                                              1 node

Activation function:                                  selu                                             linear

Batch size = 32, Number of epochs = 128, loss function = mean squared error

Optimizer: Nadam ( learning rate = 0.00001, beta_1 = 0.8, beta_2 = 0.85)

Normalized the target sine signal of NN by dividing by its maximum value.

Plot of predicted output by neural network, applied input signal & residual error given in 1st attachment. These weights of the model in the final epoch have been saved to h5 file and then loaded & tested with simulated data of 4 sine waves with amplitudes and frequencies changing with time from their initial values by random uniform noise ranging from 0 to 0.05. Plot of predicted output by neural network, target signal of sine waves & residual error given in 2nd attachment. The actual signal can be got from predicted output of NN by multiplication with normalization constant used before. However, even though network fits training  & validation sets efficiently, it gives a comparatively large error on test data of varying amplitude & frequency.

Gautam suggested to try training on this noisy data of varying amplitudes and frequencies. The results using the same model of NN is given in Attachment 3. It was found that tuning the number of nodes, layers or learning rate didn't improve fitting much in this case.

 

 

Attachment 1: nn_simulation_2_normalized_mult_sin_nodes8_128epochs_lr0p00001_beta1_0p8_beta2_0p85_0p4train_0p1valid_marked.pdf
nn_simulation_2_normalized_mult_sin_nodes8_128epochs_lr0p00001_beta1_0p8_beta2_0p85_0p4train_0p1valid_marked.pdf
Attachment 2: nn_simulation_normalizedtarget_128epochs_mult_sin_load_wt_varyingtest_nodes8_lr0p00001_beta1_0p8_beta2_0p85_0p4train_0p1valid_marked.pdf
nn_simulation_normalizedtarget_128epochs_mult_sin_load_wt_varyingtest_nodes8_lr0p00001_beta1_0p8_beta2_0p85_0p4train_0p1valid_marked.pdf
Attachment 3: nn_simulation_2_normalized_varying_mult_sin_nodes8_128epochs_lr0p00001_beta1_0p8_beta2_0p85_0p4train_0p1valid_marked.pdf
nn_simulation_2_normalized_varying_mult_sin_nodes8_128epochs_lr0p00001_beta1_0p8_beta2_0p85_0p4train_0p1valid_marked.pdf
  14096   Sat Jul 21 14:03:19 2018 KojiSummaryThermal CompensationY arm locking

Ah. With MC2 feedback, we have about 3 times smaller "optical gain" for the ASS A2L. We have same dither, same actuator, but we need only 1/3 actuation of the MC2 compared to the ETMY case.
We had to reduce the ASS spot servo from 1 to 0.3 to make is stable, so this means that the ASS is really merginally stable.

  14095   Sat Jul 21 01:14:02 2018 gautamUpdateOMCPZT Jena driver board check

[Aaron, gautam]

We did a quick check of this board today. Main takeaways:

  • There are two voltages (HV pos and HV neg) that are output from this unit.
  • Presumably, these goto different piezoelectric elements, referenced to ground. Are there any spec sheets for these describing the geometry/threshold voltages?
  • The outputs are:
    • \mathrm{HV_{+}} = 10(V_{\mathrm{DAC}}+V_{\mathrm{offset}}), \mathrm{HV_{-}} = 10(-V_{\mathrm{DAC}}+V_{\mathrm{offset}})
    • So with V_{\mathrm{offset}} = 7.5 \mathrm{V}, we expect to be able to use +/- 7.5 V of DAC range.
  • The trim pot had to be adjusted to realize V_{\mathrm{offset}} = 7.5 \mathrm{V}​.
  • I assume 150V is some kind of damage threshold of the PZT, so there is no benefit to using 10V offset voltage (as this would result in 200 V at full range DAC voltages).

With the correct V_{\mathrm{offset}} = 7.5 \mathrm{V}, we expect 0V from the DAC to result in 0 actuation on the mirror, assuming that an equal 75V goes to 2 PZTs mounted diametrically opposite on the optic. Hopefully, this means we have sufficient range to scan the input pointing into the OMC and get some sort of signal in the REFL signal (while length PZT is being scanned) which indicates a resonance. 

We plan to carve out some IFO time for this work next week.

  14094   Sat Jul 21 01:06:49 2018 gautamSummaryThermal CompensationY arm locking

I implemented this today. For now, the LSC output matrix is set to actuate on MC2 for Y arm locking. As expected, the transmission was much more stable, and the PLL control signal RMS was also reduced by factor of ~3. MC2 control signal does pick up a large (~2000 cts) DC component over a few hours, so we need to relieve this periodically.

Now that we have a workable ASS for the Y arm as well, we should be able to have more confidence in returning to the same beam spot position on the ETM and staying there during a scan using this technique.

The main improvement to be trialled next in the scanning is to improve the speed of scanning. As things stand, my script takes ~2.5 seconds per datapoint. If we can cut this in half, that'd be huge. On Wednesday night, we were extraordinarily lucky to avoid MC3 glitching, EPICS/slow machine failures, and GPIB freezes. Today, the latter reared its head. Fortunately, since I'm dumping data to file for each datapoint, this means we at least have data till the GPIB freeze.

Quote:

For future measurements, we should consider locking the IMC length to the arm cavity - this would eliminate such alignment drifts, and maybe also make the PLL control signal RMS smaller. 


Not related to this work: Terra, Sandrine, Keerthana and I cleaned up the lab a bit today, and made better cable labels. Aaron and I have to clean up the OMC area a bit. Huge thanks to Steve for taking care of our mess elsewhere in the lab!

  14093   Fri Jul 20 22:53:15 2018 KojiUpdateASSAttempt to resurrect Yarm ASS

[Koji Gautam]

We managed to realize stable ASS configuration for Yarm. The transmission of 1.06~1.07 was recovered by introducing intentional beam spot offset in the horizontal direction towards the opposite side of the elliptic reflector. The end table optics were adjusted to have the spots about the center of the mirrors, lenses, and PDs/QPDs.


Preparation

- The Y arm was manually aligned with a given input axis. The transmission was ~0.8.
- Then, TT2 was moved in yaw such that it introduced the horizontal beam shift at the end. By moving the spot to the opposite side of the reflector. The transmission ~0.95 was obtained after patient alignment work.

- Went to the end table and checked the spots. The beam was not at the center of the last 1" lens for the Trans PDs. The beam steering was adjusted to have the spot nicely going through the lens and the mirrors. This made the transmission level to be ~1.05.

- The beam centering on the Trans PD was checked and adjusted.
- The beam centering on the RF BBPD for the arm scan was checked. The spot was too big for that PD. The lens was slightly moved away from the PD to make the spot on the BBPD small. Now the PD saw the plateu when the steering was scanned (i.e. the spot is small enough).

- With the Y arm locked with MC2, the servo gain needs to be 0.012 instead of nominal 0.015 with ETMY to prevent from servo oscilating.

ASS tuning

- First of all, only the bottom 4 loops out of total 8 loops were tuned. They are the servos for the beam alignment with regard to the caivty. The linearity and the zero crossings were checked with regard to the reference alignment. All of these 4 showed offsets that causes the servo running away. Don't know the reason of this offset, but it is freq dependent. Therefore the dither freqs were tuned to make the offset zeroed, and tuned the demod phases there. This kept the transmission as high as the reference (~1.05)

- This allowed us to play with the spot position a bit by tuning the caivty alignment. In the end, the transmission of ~1.08 was obtained. Using this alignment, A2L offset for ETMY Yaw was determined to be +17 (to make the error signal -17). This offset produces almost a beam radius (5mm) shifted on the end mirror towards the opposite direction of the reflector.

- The nominal servo setting made the spot servo running away. Gautam pointed out that this could be a gain hierarchy problem (i.e. the spot servos are too fast). We ended up reducing the gain of the servo from 1.0 to 0.3 to make the spot servo stable.

- All the ASS setting was stored in a new snap file "script/ASS/ASS-DITEHR_ON.snap". The previous snap was saved to "script/ASS/ASS_DITHER_ON_preVent201807.snap". This did not save the exc gains of the oscillators. Therefore "DITHER_ASS_ON.py" was modified to have the new exc gains (CLKGAIN). The old values are stored in the comments in this script.


Overall this is not an ideal situation as we don't know what is the actually cause of the offsets in the dither error signals. We expect to correct the beam clipping and the suspension sooner or later. Therefore, we will come back to the ASS again once the other issues are corrected.

Attachment 1: 02.png
02.png
  14092   Fri Jul 20 22:51:28 2018 KojiUpdateIOOIMC WFS path alignment

IMC WFS tuning

- IMC was aligned manually to have maximum output and also spot at the center of the end QPD.
- The IMC WFS spots were aligned to be the center of the WFS QPDs.
- With the good alignment, WFS RF offset and MC2 QPD offsets were tuned via the scripts.

  14091   Fri Jul 20 18:30:47 2018 JonConfigurationAUXRecommend to install AUX PZT driver

I recently realized that the PLL is only using about 20% of the available actuation range of the AUX PZT. The +/-10 V control signal from the LB1005 is being directly inputted into the fast AUX PZT channel, which has an input range of +/-50 V.

I recommend to install a PZT driver (amplifier) between the controller and laser to use the full available actuator range. For cavity scans, this will increase the available sweep range from +/-50 MHz to +/-250MHz. This has a unique advantage even if slow temperature feedback is also implemented. To sample faster than the timescale of most of the angular noise,  scans generally need to be made with a total sweep time <1 sec. This is faster than the PLL offset can be offloaded via the slow temperature control, so the only way to scan more than 100 MHz in one measurement is with a larger dynamic range.

  14090   Fri Jul 20 07:43:54 2018 SteveSummarySUSETMY

 

 

Attachment 1: ETMY_leveling.png
ETMY_leveling.png
Attachment 2: ETMY.png
ETMY.png
  14089   Thu Jul 19 18:09:17 2018 poojaUpdateCamerasUpdate in developing neural networks

Aim: To develop a neural network that resolves mirror motion from video.

Case 1:

Input : Simulated video of beam spot motion in pitch by applying 4 sine  waves of frquencies 0.2, 0.4, 0.1, 0.3 Hz  and amplitude ratios to frame size to be 0.1, 0.04, 0.05, 0.08

The data has been split into train, validation and test datasets and I tried training on neural network with the same model topology & parameters as in my previous elog (https://nodus.ligo.caltech.edu:8081/40m/14070)

The output of NN and residual error have been shown in Attachment 1. This NN model gives a large error for this. So I think we have to increase the number of nodes and learning rate so that we get a lower error value with a single sine wave simulated video ( but not overfitting) and then try training on linear combination of sine waves.

Case 2 :

Normalized the target sine signal of NN so that it ranges from -1 to 1 and then trained on the same neural network as in my previous elog with simulated video created using single sine wave. This gave comparatively lower error (shown in Attachment 2). But if we train using this network, we can get only the frequency of test mass motion but we can't resolve the amount by which test mass moves. So I'm unclear about whether we can use this.

Attachment 1: nn_simulation_mlt_sine_nodes4_lr0p00001_beta1_0p8_beta2_0p85_marked.pdf
nn_simulation_mlt_sine_nodes4_lr0p00001_beta1_0p8_beta2_0p85_marked.pdf
Attachment 2: nn_simulation_2_nodes4_target-1to1_marked.pdf
nn_simulation_2_nodes4_target-1to1_marked.pdf
  14088   Thu Jul 19 13:35:30 2018 SteveSummaryVACannuloses pumped

Roughing down the annuloses required closing V1 for 13 minutes

IFO is 2.2e-5 Torr

Attachment 1: AnsPumped.png
AnsPumped.png
  14087   Thu Jul 19 11:01:03 2018 SteveSummaryVACpd81 @ 2e-5 Torr

Cold cathode gauge just turned on.

Attachment 1: pd81@2days.png
pd81@2days.png
  14086   Thu Jul 19 04:44:09 2018 Annalisa, TerraSummaryThermal Compensationfrequency shift observed with heating!

Annalisa, Gautum, Koji, Terra

Summary: with the reflector setup, we measured a frequency shift of the first and second order modes! First looks of shifts show 1st HOM shift ~-10 kHz, 2nd HOM shift ~-18 kHz (carrier ~4 kHz). We saw no shift with the cylinder/lenses set up.

- - - - -

Tonight we modified the cavity scan setup: the LO is provided by the Marconi which, at the same time, is also used to scan the AUX laser frequency instead of the Agilent. In order to get rid of the free running noise between Marconi and Agilent, the Marconi frequency was scanned and, point by point, the Agilent center frequency was changed accordingly. In order to speed up the process, the whole procedure was automated. The script is called AGfast.py and can be found in /users/annalisa/postVent.

One thing that helped in improving the data quality of the phase information was to set the Agilent IF bandwidth @1kHz. Not yet clear why, but it was better than having a lower bandwidth. To be further investigated.

With this setup, we made some coarse scan of the full FSR and then we "zoomed" around the main peaks in order to increase the resolution and get a more precise information about the peak frequency.

Here are the frequency ranges that we scanned:

  • carrier - central frequency: 31.73MHz; range: [31.68MHz - 31.78MHz]
  • HOM1 - central frequency: 32.88MHz; range: [32.84MHz - 32.93MHz]
  • HOM2 - central frequency: 34.03MHz; range: [33.95MHz - 34.06MHz]
  • HOM3 - central frequency: 35.18MHz; range: [35.09MHz - 35.25MHz]

We powered the heater of the lenses setup @4:55 am at 14.4V and 0.9A. Then we slightly increased the power @5:05am and the final "hot state" configuration is with heater powered at 16V and 0.9A.

With this setup we couldn't see any frequency shift

Then, at around 6:30 am we turned on the reflector setup and we measured a frequency shift of the first and second order modes. First scans show 1st HOM shift ~10 kHz, 2nd HOM shift ~18 kHz. First attachment shows carrier hot/cold, second attachment shows HOM2 hot/cold. We started to get plauged by high seismic noise. Heaters turned off at 7:45 am. Lots of scans and actual analysis to go.


gautam: about the questionable plotting -

  • 10 faint (alpha~0.3) lines are individual measurements with the reflector doing its heating. (AG4395A, 0 span, single frequency measurements plotted together).
  • charcoal line, labelled mean, is the mean of the 10 above lines.
  • bright green ("Reference") is the mean of a coarse scan (cold ETM) overlaid for comparison. 
  • "cold" - self explanatory.

My personal favourite plot is Attachment #3, which is a 5 MHz scan (cold) to identify positions of the various peaks. The power of including phase information in the analysis is clear. The second FSR on the right edge of the plot is not as prominent as the first is because the arm transmission was degrading throughout the measurement. For future measurements, we should consider locking the IMC length to the arm cavity - this would eliminate such alignment drifts, and maybe also make the PLL control signal RMS smaller. 

Attachment 1: scanning_fine_2018-07-19-07-32-08_parsed.pdf
scanning_fine_2018-07-19-07-32-08_parsed.pdf
Attachment 2: scanning_fine_2018-07-19-06-57-47_parsed.pdf
scanning_fine_2018-07-19-06-57-47_parsed.pdf
Attachment 3: Yscan_scanning_parsed_2am.txt.pdf
Yscan_scanning_parsed_2am.txt.pdf
  14085   Thu Jul 19 01:56:25 2018 gautamSummaryVACAUX pump shutdown

[koji, gautam]

Per Steve's instructions, we did the following:

  • TP3fl pressure reading was 65 torr.
  • TP3 controller reported pumping current of ~0.18A, temperature of 24C.
  • We throttled the manual valve which was connecting the "AUX" pump to the TP3fl.
  • The TP3fl pressure went up to 330 torr.
  • TP3fl controller reported current of 0.22A, temperature of 24C.
  • After ~5mins, we shut the AUX pump off.
  • We have monitored it over the last 1hour, no red flags.
    • (Before stopping AUX RP)
      0:56AM TP3 I=0.18A, P=6W, 23degC, TP3FL: 66
    • 0:59AM TP3 I=0.22A, P=7W, 23degC, TP3FL: 336
    • 1:15AM TP3 I=0.21A, P=7W, 23degC, TP3FL: 320
    • 1:31AM TP3 I=0.21A, P=7W, 23degC, TP3FL: 310
    • 2:06AM TP3 I=0.21A, P=7W, 23degC, TP3FL: 301
    • 5:06AM TP3 I=0.21A, P=7W, 23degC, TP3FL: 275
  14084   Wed Jul 18 23:43:50 2018 KojiUpdateGeneralVent 80 recovery

Is the reflector too close to the beam and causing clipping?

Quote:

For unknown reasons, the Y arm ASS does not maximize TRY. So we are in the unfortunate situation of neither arm having a working ASS servo. To be worked on later.

Attachment 1: IMG_5868.JPG
IMG_5868.JPG
Attachment 2: IMG_5382.JPG
IMG_5382.JPG
  14083   Wed Jul 18 17:36:50 2018 SteveSummaryVACpumpdown 81 at 6 +9 hrs completed

IFO P1 6e-4 Torr,  manual gate valve is fully open

The annuloses will be pumped down tomorrow.

Valve configuration: vacuum normal, RGA and annuloses are not pumped

Quote:

The manual gate valve scan was clean. Atm1     TP1 was pumping on it overnight.

                                                Pumpdown continued to hand over the pumping to TP1 maglev turbo

V1 was opened at P1 400 mTorr  with manual gate at 3/4 turn open position as Magev at 560 Hz rotation.

This is the first time we pumping down from atm with one small "beer can" turbo  and throttled gate to control load on small turbo forepump

The 70 l/s turbo is operating at 50k RPM, 0.7 A and 31 C,  pumping speed  ~ 44 mTorr/h at 200-400 mTorr range.

Watching foreline pressures and current one can keep opening gate valve little by little the so the load is optimized. It is working but not fast.

Let's keep small turbo at 0.8 Amp and 32 C max at this pumpdown. 

Quote:

10:20PM

  • Opened VM2 to pump down the RGA section with TP1
  • Stopped rotary roughing pumps
    • Manually closed RV1
    • Closed V3
    • Stopped RP1 and RP3
    • Vented the RP hose

The P1 pressure is 380mTorr. I allowed Gautam to use the full PSL power (~1W).

Attachment 1: pd81completed.png
pd81completed.png
Attachment 2: pd81@30hrs.png
pd81@30hrs.png
  14082   Wed Jul 18 12:49:08 2018 SteveSummaryVACpumpdown 81 at 6 +4.5hrs

The manual gate valve scan was clean. Atm1     TP1 was pumping on it overnight.

                                                Pumpdown continued to hand over the pumping to TP1 maglev turbo

V1 was opened at P1 400 mTorr  with manual gate at 3/4 turn open position as Magev at 560 Hz rotation.

Two aux fans on to hold tubo temps TP1 & TP3 . Atm3

This is the first time we pumping down from atm with ONE small "beer can" turbo  and throttled gate valve to control load on small turbo forepump

The 70 l/s turbo is operating at 50k RPM, 0.7 A and 31 C,  pumping speed  ~ 44 mTorr/h at 200-400 mTorr range with aux drypump in the foreline of TP3

Watching foreline pressures and current one can keep opening gate valve little by little the so the load is optimized. It is working but not fast.

Let's keep small turbo at 0.8 Amp and 32 C max at this pumpdown. 

Quote:

10:20PM

  • Opened VM2 to pump down the RGA section with TP1
  • Stopped rotary roughing pumps
    • Manually closed RV1
    • Closed V3
    • Stopped RP1 and RP3
    • Vented the RP hose

The P1 pressure is 380mTorr. I allowed Gautam to use the full PSL power (~1W).

Attachment 1: manlGateScan.png
manlGateScan.png
Attachment 2: handing_over_Mag.png
handing_over_Mag.png
Attachment 3: TGVw2auxfans_.jpg
TGVw2auxfans_.jpg
  14081   Wed Jul 18 03:14:48 2018 AnnalisaUpdateGeneralVent 80 recovery

[Gautam, Johannes, Koji, Annalisa]

Tonight we increased the power of the PSL laser and we achieved the lock of both arms with high power.

The AUX beam alignment to the Y arm was recovered and the PLL restored (using the Marconi as LO).

We made a quick measurement of the phase noise and the results will be posted tomorrow.

The beam on the PSL has been blocked, as well as the AUX beam on the AS table. The Marconi has been switched off.


gautam:

  1. Before turning up PSL power, I placed a block in front of MC refl to avoid any PD burning. Replaced HR Y1 2" optic with the usual 10% reflective BS to direct MC REFL to the locking PD.
  2. Waveplate was rotated back to 180 deg (original position before the vent). After optimizing PMC transmission, I measured 1.05 W going into the IMC (pre-vent value was 1.07 W, prolly within power meter absolute accuracy).
  3. IMC autolocker restored to usual high power version on megatron.
  4. There seems to be some kind of vacuum interlock in effect that prevents me from opening the PSL shutter via EPICS - I had to toggle the position on the shutter controller under the table. After tonight's work, I returned the controller to the NC state, to avoid any further interference with this interlock code that may prevent pumping in the AM.
  5. PLL gain was re-adjusted to achieve maximum stability (judged by eye) of the beat-note in lock triggered on the Marconi LO signal. Alignment onto the NF beatPD was also tweaked to squeeze out as much beat as possible.
  6. The main objective tonight was to send AUX beam in, recover transmission beat, scan the AUX frequency, and resolve some peaks (MAX HOLD scanning technique, magnitude only for now, no phase info). Thanks to JE's expert fiber alignment and beatnote maximization, we achieved this yes. Annalisa will post a plot tmr. 
  7. For unknown reasons, the Y arm ASS does not maximize TRY. So we are in the unfortunate situation of neither arm having a working ASS servo. To be worked on later.
  14080   Tue Jul 17 22:25:41 2018 KojiSummaryVACpumpdown 81 at 6 hrs

10:20PM

  • Opened VM2 to pump down the RGA section with TP1
  • Stopped rotary roughing pumps
    • Manually closed RV1
    • Closed V3
    • Stopped RP1 and RP3
    • Vented the RP hose

The P1 pressure is 380mTorr. I allowed Gautam to use the full PSL power (~1W).

  14079   Tue Jul 17 18:16:38 2018 SteveSummaryVACpumpdown 81 at 6 hrs

Precondition:  4 days at atm.   Atm5

HEPA tent used during the vent at ETMY  It reduced partical count 10 fold of 0.5 and 0.3 micron particals. Atm6

New items in vacuum:  Clean manual gate valve [Cetec made] from John Worden with 6" id....as it came from Hanford... [ Throttle able gate valve- TGV ] Atm3

                                 ( note: we have 3 more identical in the lab. The original intention was to use them for purging gates )

                                  Optiform Au plated reflector , Induceramics heating elements, similar as existing Cooner cables and related lenses, hardwear. see 14078

                                  OMC related item : none......... 14,110

 

The pumpdown is at 510 mTorr with RP1 & RP3 still pumping. Koji will shut it down the roughing later tonight. Tomorrow morning I will start the pumping by switching over to TP1 maglev.

Thanks for Koji and Gautam'  help of the installation of the manual gate valve. Atm4  This will allow us to control the load on our Varian foreline D70 turbo TP3

 

Attachment 1: pd81@6hrs.png
pd81@6hrs.png
Attachment 2: before_c.jpg
before_c.jpg
Attachment 3: tgv_c.jpg
tgv_c.jpg
Attachment 4: TGVinstalled.jpg
TGVinstalled.jpg
Attachment 5: 4_days_vent.png
4_days_vent.png
Attachment 6: tentHEPA.jpg
tentHEPA.jpg
  14078   Tue Jul 17 17:37:46 2018 Annalisa, TerraConfigurationThermal CompensationHeaters installation

Summary

We installed two heaters setup on the ETMY bench in order to try inducing some radius of curvature change and therefore HOMs frequency shift.

------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------

We installed two heaters setup.

Elliptic reflector setup (H1): heater put in the focus of the elliptical reflector: this will make a heat pattern as descirbed in the elogs #14043 and #14050. 

 

Lenses setup (H2): heater put in a cylndrical reflector (made up with aluminum foil) 1'' diameter, and 2 ZnSe lenses telescope, composed by a 1.5'' and a 1'' diameter respectively, both 3.5'' focal length. The telescope is designed in such a way to focus the heat map on the mirror HR surface. For this latter the schematic was supposed to be the following: 

This setup will project on the mirror a heat pattern like this:

which is very convenient if we want to see a different radius of curvature for different HOMs. However, the power that we are supposed to have absorbed by the mirror with this setup is very low (order of 40-ish mW with 18V, 1.2A) which is probably not enough to see an effect. Moreover, mostly for space reasons (post base too big), the distances were not fully kept, and we ended up with the following setup:

In this configuration we won't probably have a perfect focusing of the heat pattern on the mirror.

In vacuum connections

See Koji's elog #14077 for the final pin connection details. In summary, in vacuum the pins are:

13 to 8 --> cable bunch 0

7 to 2 --> cable bunch 2

25 to 20 --> cable bunch 1

19 to 14 --> cable bunch 3

where Elliptic reflector setup (H1) is connected to cables 0 and 1, and the lenses setup is connected to cables 2 and 3.

Installed setup

This is the installed setup as seen from above:

Attachment 5: IMG_5380.JPG
IMG_5380.JPG
  14077   Tue Jul 17 12:55:45 2018 KojiSummaryGeneralStarted pumping

[Steve, Koji, Gautam]

We started pumping down at ~12:15PM.

Vent finalization ~ YEND

  • The table leveling was way off. This was adjusted by the balancing weight. (Attachment 1~3)
  • The alignment of ETMY was not too much off. Just aligned it with the oplev spot on MEDM and this already made the green flashing.
  • The Green TEM00 was maximized with ITMY and ETMY. This made the PSL IR flashing.
  • The heater wires were checked. I found that one of the heater wires was touching the optical table via the cable shield. This is because the upper pins were shifted to the left side (Attachment 4&5). The pins were shifted and now all 4 cables are isolated form the table. I also checked the mutual resistance between the 4 terminals. They were measured to be isolated except two pairs that showed 4.4 Ohms and 4.0 Ohms (Attachment 6)
  • The tools were removed from the chamber. The Y arm was still flashing.
  • We closed the ETMY door.

Vent finalization ~ Vertex

  • Found the ITMX stuck. Gautam came in and showed us his black magic to release the optic...
  • This allowed us to align X arm. The green flash was found and the TEM00 flash was seen. This allowed us to see the PSL IR flash at the X end.
  • PRM Refl was aligned. SRM was aligned with the oplev.
  • The beam on the AS port was checked. The AS beam came out from the window.
  • Closed the OMC chamber.

Pumping

  • Started pumping with RP1 and RP3. (~12:15PM)

Attachment 1: IMG_5408.JPG
IMG_5408.JPG
Attachment 2: IMG_5400.JPG
IMG_5400.JPG
Attachment 3: IMG_5401.JPG
IMG_5401.JPG
Attachment 4: IMG_5402.JPG
IMG_5402.JPG
Attachment 5: IMG_5403.JPG
IMG_5403.JPG
Attachment 6: IMG_5404.JPG
IMG_5404.JPG
  14076   Tue Jul 17 12:46:28 2018 ranaUpdateGeneralsome notes from yesterday

For the EY, instead of balancing the table, I just moved the weight approximately so that the ETMY OSEMS were at half light, but didn't check the level since ETMY is the only optic.

 

Some notes on OMC/AS work (Aaron/Gautam can amend/correct):

- Beam is now well centered in OMC MMT. Hits input coupling mirror and cleanly exits the vacuum to the AS table.

- Didn't see much on OMC trans, but PDs are good based on flashlight test.

- just before closing, re-aligned beam in yaw so that it gets close to the east screw on the input coupler. Aaron and I think we maybe saw a flash there with the OMC length PZT being driven at full range by a triangle wave.

- with OMC Undulators (aka tip/tilt PZT mirrors) energized, the beam was low on PZT1 mirror. We pitched ITMY by ~150 micro-rad and that centered the beam on PZT1 mirror. ITMY-OL is probably not better than 100 urad as a DC reference?

- We checked the range of Undulator 1 and we were getting ~5 mrad of yaw of the beam for the full range, and perhaps half of that in pitch. Rob Ward emailed us from Oz to say that the range is robably 2.7 mrad, so that checks out.

Even if the ITMY has to be in the wrong position to get the beam to the OMC, we can still do the heater tests in one position and then do the OMC checkout stuff in the other position.

Gautam suspects that there is a possible hysterical behaviour in the Undulators which is related to the MC3 glitching and the slow machine hangups and also possibly the illuminati.

[aaron]

-We noticed a ghost beam that from MC REFL (MMT2) that should be dumped during the next vent--it travels parallel to the OMC's long axis and nearly hits one of the steering mirrors for OMC refl.

-We measured the level of the table and found it ~3 divisions off from level, with the south end tilted up

-Gautam rotated and slightly translated OM5 to realign the optic, as expected. No additional optics were added.

-Gautam and I tested the TT piezo driver. We found that 3.6V on the driver's input gave 75V (of 150V) at the output, at least for yaw on piezo 1. However, as Gautam mentioned, during testing it seemed that the other outputs may have different (nonzero) offset voltages, or some hysterisis.

  14075   Tue Jul 17 01:07:40 2018 gautamUpdateSUSETMY EQ stops

For the heater setup on EY table, I EQ-stopped ETMY. Only the face EQ stops (3 on HR face, 2 on AR face) were engaged. The EY Oplev HeNe was also shutdown during this procedure. 

  14074   Mon Jul 16 18:12:00 2018 KojiUpdateVACAdding a manual gate valve between TP1 and V1/VM2

[Steve Koji]

We are in the process of adding a manual gate valve between TP1 (Osaka Maglev) and the other gate valves (I suppose V1 and VM2).
The work is still on going and we will continue to work on this tomorrow. Because this section is isolated from the main volume, this work does not hold off the possible rough pumping tomorrow morning.

The motivation of this work is as follows:
- Since TP2 failed, the main vacuum volume has been pumped down by TP1 and TP3. However TP3 is not capable to handle the large pressure difference at the early stage of the turbo pumping. This cause TP3 to have excessive heating or even thermal shutdown.
- The remedy is to put a gate valve between TPs and the main vacuum to limit the amount of gas flowing into the TPs. This indeed slows down the pumping speed of turbo, but this is not the dominant part of the pumping time.

Actual work:
- Comfirmed TP1 is isolated.
- Unscrewed the flange of TP1.
- Remove TP1. This required to lift up TP1 with some shim as the nuts interferes with the TP1 body. (Attachment1, 2, 3)
- Now remove 10inch flange adapter. (Attachment4)
-
Attach 10"-8" adapter and 8" rotational sleeve. (Attachment5)

Attachment 1: P_20180716_155413.jpg
P_20180716_155413.jpg
Attachment 2: P_20180716_155645.jpg
P_20180716_155645.jpg
Attachment 3: P_20180716_155738.jpg
P_20180716_155738.jpg
Attachment 4: P_20180716_162307.jpg
P_20180716_162307.jpg
Attachment 5: P_20180716_172000.jpg
P_20180716_172000.jpg
  14073   Mon Jul 16 15:07:19 2018 KojiSummaryVACOven C vent

[Steve Koji]

- Attachment1: Removed the thermal cap. Checked the temperature of the oven. It was totally cold.

- Attachment2: Confirmed the RGA section was isolated. The pumps for the RGA was left running.

- Attachment3: Closed the main valve. The pumps for the main volume was left running.

- Attachment4: Started removing the rid. This did not change the gause readings as they were isolated from the venting main volume.

- Attachment5: Opened the rid. Took the components out on a UHV foil bag. The rid was replaced but loosely held by a few screws with the old gasket, just to protect the frange and the volume from rough dusts.

Attachment 1: P_20180716_141512.jpg
P_20180716_141512.jpg
Attachment 2: P_20180716_141601.jpg
P_20180716_141601.jpg
Attachment 3: P_20180716_141610.jpg
P_20180716_141610.jpg
Attachment 4: P_20180716_141827.jpg
P_20180716_141827.jpg
Attachment 5: P_20180716_143901.jpg
P_20180716_143901.jpg
  14072   Sat Jul 14 16:04:34 2018 aaronUpdateOMCChecking OMC Electronics

Next check is the DCPD/OMMT Satellite Box

I traced a cable from the OMC electrical feedthrough flanges to find the DCPD/OMMT Satellite Box (D060105). I couldn't find the DCC number or mention of the box anywhere except this old elog.

Gautam and I supplied the box with power and tested what we think is the bias for the PD, but don't read any bias... we tracked down the problem to a suspicious cable, labelled.

We confirmed that the board supplies the +5V bias that Rich told us we should supply to the PDs.

We tested the TFs for the board from the PD input pins to output pins with a 100kHz low pass (attached, sorry no phase plots). The TFs look flat as expected. The unfiltered outputs of the board appear bandpassed; we couldn't identify why this was from the circuit diagram but didn't worry too much about it, as we can plan to use the low passed outputs.

Attachment 1: Screenshot_2018-07-14_17.53.40.png
Screenshot_2018-07-14_17.53.40.png
Attachment 2: Screenshot_2018-07-14_17.57.17.png
Screenshot_2018-07-14_17.57.17.png
  14071   Fri Jul 13 23:39:46 2018 AnnalisaConfigurationThermal CompensationThermal compensation setup - power supply

[Annalisa, Rana]

In order to power the heater setup to be installed in the ETMY chamber, we took the Sorensen DSC33-33E power supply from the Xend rack which was supposed to power the heater for the seismometer setup.

We modified the J3 connector behind in such a way to allow a remote control (unsoldered pins 9 and 8). 

Now pins 9 and 12 need to be connected to a BNC cable running to the EPICS.


RXA update: the Sorensen's have the capability to be controlled by an external current source, voltage source, or resistive load. We have configured it so that 0-5V moves the output from 0-33 V. There is also the possibility to make it a current source and have the output current (rather than voltage) follow the control voltage. This might be useful since out heater resistance is changing with temperature.

Attachment 1: IMG_2012.jpg
IMG_2012.jpg
Attachment 2: IMG_2013.jpg
IMG_2013.jpg
Attachment 3: 20180713_213818.jpg
20180713_213818.jpg
  14070   Fri Jul 13 23:23:49 2018 poojaUpdateCamerasUpdate in developing neural networks

Aim: To develop a neural network that resolves mirror motion from video.

I tried to reduce the overfitting problem in previous neural network by reducing the number of nodes and layers and by varying the learning rate, beta factors (exponential decay rates of moving first and second moments) of Nadam optimizer assuming error of 5% is reasonable.

Input:

32 * 32 image frames (converted to 1d array & pixel values of 0 to 255 normalized) of simulated video by applying sine signal to move beam spot in pitch with frequency 0.2Hz and at 10 frames per second.

Total: 300 cycles ,           Train: 60 cycles,    Validation: 90 cycles,    Test: 150 cycles

Model topology:

                                          Input               -->                  Hidden layer               -->                    Output layer                                  

                                                                                          4 nodes                                              1 node

Activation function:                                  selu                                             linear

Batch size = 32, Number of epochs = 128, loss function = mean squared error

Optimizer: Nadam

Case 1:

Learning rate = 0.00001,    beta_1 = 0.8 (default value in Keras = 0.9),  beta_2 = 0.85 (default value in Keras = 0.999)

Plot of predicted output by neural network, applied input signal & residual error given in 1st attachment.

Case 2:

Changed number of nodes in hidden layer from 4 to 8. All other parameters same.

These plots show that when residual error increases basically the output of neural network has a smaller amplitude compared to the applied signal. This kind of training error is unclear to me.

When beta parameters of optimizer is changed farther from 1, error increases.

Attachment 1: nn_simulation_2_nodes4_lr0p00001_beta1_0p8_beta2_0p85.pdf
nn_simulation_2_nodes4_lr0p00001_beta1_0p8_beta2_0p85.pdf
Attachment 2: nn_simulation_2_nodes8_lr0p00001_beta1_0p8_beta2_0p85.pdf
nn_simulation_2_nodes8_lr0p00001_beta1_0p8_beta2_0p85.pdf
  14069   Fri Jul 13 20:36:33 2018 KojiSummaryGeneralIn vac/In air heater wiring

I went to the Y-end and took more photos of the cable stand. These revealed that in-vac pin #13 is connected to the shield of the cable (P.2). This in-vac pin #13 corresponds to  in-air pin #1. So in the end, we bunch the pins in the following order.

In Air In Vac
Pin #2-7 Pin #12-7
Pin #8-13 Pin #6-1
Pin #14-19 Pin #25-20
Pin #20-25 Pin #19-14

 

Attachment 1: heater_wiring.pdf
heater_wiring.pdf heater_wiring.pdf heater_wiring.pdf
  14068   Fri Jul 13 18:01:13 2018 gautamUpdateGeneralLow power MC

After getting the go ahead from Steve, I removed the physical beam block on the PSL table, sent the beam into the IFO, and re-aligned the MC to lock at low power. I've also revived my low power autolocker (running on megatron), seems to work okay though the gains may not be optimal, but it seems to do the job for now. Nominal transmission when well aligned at low power is ~1200cts. I briefly checked Y arm alignment with the green, seems okay, but didn't try locking the Y arm yet. All doors are still on, and I'm closing the PSL shutter again while Keerthana and Sandrine are working near the AS table.

ELOG V3.1.3-